OpenStudy (anonymous):

Solve for x: 4 over 5 x + 4 over 3 = 2x

OpenStudy (anonymous):

Hello! So is this the problem you are talking about? Four fifths X plus four thirds equals 2x?

OpenStudy (anonymous):

aja

OpenStudy (happy_to_help):

\[\frac{ 5(4)+4 }{ 3=2(4) }\]

OpenStudy (happy_to_help):

First multiply 5x4=20 20+4=24 then multiply 2x4=8 3=8